If 3.2443 moles of carbon dioxide are produced, how many grams of iron (III) oxalate were reacted if the reaction proceeds with a 99.6% yield?



1
Expert's answer
2022-04-19T03:11:54-0400

Solution:

carbon dioxide = CO2

iron(III) oxalate = Fe2(C2O4)3


Assume that the reaction is the decomposition of iron(III) oxalate

Then, the balanced chemical equation is:

Fe2(C2O4)3 → 2FeC2O4 + 2CO2

According to stoichiometry:

1 mol of Fe2(C2O4)3 produces 2 mol of CO2

X mol of Fe2(C2O4)3 produces 3.2443 mol of CO2

Thus,

Moles of Fe2(C2O4)3 = X = (3.2443 mol CO2) × (1 mol Fe2(C2O4)3 / 2 mol CO2) = 1.62215 mol Fe2(C2O4)3

Given the yield of the reaction:

Moles of Fe2(C2O4)3 = (1.62215 mol × 100%) / (99.6%) = 1.628665 mol


The molar mass of Fe2(C2O4)3 is 375.747 g mol−1

Therefore,

Mass of Fe2(C2O4)3 = (1.628665 mol) × (375.747 g mol−1) = 611.966 g

Mass of Fe2(C2O4)3 = 611.966 g


Answer: 611.966 grams of iron(III) oxalate were reacted